Internet providers: In a survey of

686 homeowners with high-speed Internet, the average monthly cost of a high-speed Internet plan was $54.86 with standard deviation $12.1. Assume the plan costs to be approximately bell-shaped. Estimate the number of plans that cost between $42.76 and $66.96. Round to the nearest whole number.

The number of plans that cost between $42.76 and $66.96 is
Answer #1

Solution :

Given that,

Using Empirical rule,

P( - 1< X < + 1) = 68%

P(54.86 - 12.1 < X < 54.86 + 12.2) = 68%

P(42.76 < X < 66.96) = 68%

Answer = 686 * 68% = 466.48 = 466